OpenStudy (anonymous):

Magma that does not reach the surface is called: intrusive rock extrusive rock metamorphic rock imbricated rock

OpenStudy (anonymous):

can you help me ???

OpenStudy (anonymous):

hey back again !!!

OpenStudy (study312):

it's A.

OpenStudy (anonymous):

Igneous rocks are made mainly of compounds of: carbon and hydrogen silicon and oxygen carbon and silicon carbon and oxygen

OpenStudy (study312):

If you need me you can tag me. by typing @ then my username

OpenStudy (anonymous):

@Study312

OpenStudy (anonymous):

like that

OpenStudy (study312):

yes

OpenStudy (anonymous):

wow this is awesome

OpenStudy (study312):

It's B

OpenStudy (anonymous):

now to my last question

OpenStudy (study312):

I'll try to tell you why it's the answer when I can

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

finally moving on to math

OpenStudy (anonymous):

my last subject of the day

OpenStudy (study312):

ok

OpenStudy (anonymous):

Which of the following is a true statement? The opposite of −45 is equal to the absolute value of −45. The opposite of 45 is equal to the absolute value of 45. The opposite of 45 is equal to the absolute value of −45. The opposite of −45 is not equal to the absolute value of 45.

OpenStudy (study312):

[-45] = 45 the absolute vaule of -45 is 45. in other words -45 is the opposite of 45 and 45 is the opposite of -45

OpenStudy (study312):

Cool beans?

OpenStudy (anonymous):

no frozen beans lol

OpenStudy (study312):

haha. so that's it?

OpenStudy (anonymous):

so its c or what

OpenStudy (study312):

oh right. Yeah it's C

OpenStudy (anonymous):

Which of the following statements is false? -|-5| = -5 -(-5) = 5 -|5| = -5 |-5| = -5

OpenStudy (anonymous):

thanx for the metal tho

OpenStudy (study312):

hmm this is a tricky one

OpenStudy (anonymous):

i think its a

OpenStudy (study312):

Let me ask someone and see what they think, one sec

OpenStudy (anonymous):

k

OpenStudy (study312):

Ok @Ms-Brains what do you think about the question Which of the following statements is false? -|-5| = -5 -(-5) = 5 -|5| = -5 |-5| = -5

MsBrains (ms-brains):

Alright gimme a sec. Let me see

OpenStudy (anonymous):

???

OpenStudy (study312):

ok Thx for coming

OpenStudy (study312):

@superkiller54 Ms-Brains has helped me in the past so I asked her to help

OpenStudy (anonymous):

k

OpenStudy (anonymous):

where is she tho ???

OpenStudy (study312):

idk

OpenStudy (study312):

my computer was freezing for 5 min

OpenStudy (anonymous):

same here

OpenStudy (study312):

maybe hers did the same

OpenStudy (study312):

her's*

OpenStudy (anonymous):

probably

OpenStudy (study312):

brb

OpenStudy (study312):

now she is offline so I think her computer did freeze

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Which of the following statements is false? If a number is a natural number, then it is rational. If a number is a whole number, then it is rational. If a number is a fraction, then it is rational. If a number is an integer, then it is irrational.

OpenStudy (anonymous):

in the meantime

OpenStudy (study312):

one sec